OpenStudy (anonymous):

A bag contains three red, seven yellow and 5 blue marbles. If three balls are drawn from the bag simultaneously , fin the probability that: (a) all balls are of different colours

OpenStudy (anonymous):

total amount of balls: 15

OpenStudy (anonymous):

we need to choose 3 different colors, so, we can do 15 C 3, i believe

OpenStudy (anonymous):

15*14*13 --------- 3*2*1

OpenStudy (anonymous):

well the answer is 3/13

OpenStudy (anonymous):

isn't 15c 3 the sample space

OpenStudy (zarkon):

no...it is the number of elements in the sample space

OpenStudy (anonymous):

yeah i know, we need to find the probability, but im a little rusty on stats, maybe somebody well have pity on me an help me out

OpenStudy (zarkon):

\[\frac{{3\choose 1}{7\choose 1}{5\choose 1}}{{15\choose 3}}\]

OpenStudy (zarkon):

\[=\frac{3\times7\times5}{455}=\frac{3}{13}\]

OpenStudy (anonymous):

oh i see thanks you very much zarkon:)